LControl

LANC/Control-L Interface for Amiga Computers

Version 1.01 Mai 1994

by Juergen Frank und Michael Watzl

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1993-94 Juergen Frank und Michael Watzl

1 What is LControl

LControl is a program to control video cameras via a special control hardware. The cameras must have a control-L or LANC interface (Sony kompatible).

The hardware is very easy to construct. Schematics, construction kit or the complete hardware will be sendt to you after registering, See Appendix B [Registering], page 7.

The control hardware is connected to the parralel port. Two cameras can be connected to the hardware. You can switch between the two cameras via the software.

LControl is a commodity and it has an arexx port. Therefore LControl can work in the background and your camera can be "programmed" via arexx.

See also Appendix C [IR-Master-Support], page 9.

2 Installation

Requirements:

• Kickstart 2.04 or higher

• Workbench 2.1 or higher for using locale. [version 1.2 of LControl] The installation of LControl is really easy. There is nothing to do! You need no additional

libraries, no assigns must be done.

When you drag LControl to the ‘WBStartup’-drawer, you should also add the DONOTWAIT tooltype.

3 How To Use

3.1 Menus

3.1.1 Project

About... Some information about the authors and the program Hide Closes the LControl window. However LControl stays active. Fix position

Snapshots the current position. This piece of information is saved in the tooltypes, See Section 3.4 [Tooltypes], page 5.

Quit Quits LControl

3.1.2 Special

Reset counter

Resets the counter of the camera.

Swap cameras

Swaps to the other camera.

Zoom in Sends "Zoom In" to the camera Zoom out Sends "Zoom Out" to the camera Record Sends "Record" to the camera.

3.2 Gadgets

Play Sends a "Play"-command to the camera Stop Sends a "Stop"-command to the camera Pause Sends a "Pause"-command to the camera FREW Sends a "fast rewind" command to the camera FFOR Sends a "fast forward" to the camera PREW Sends a "picture rewind" to the camera PFOR Sends a "picture forward" to the camera −− Sends a "slow motion rewind" to the camera − Sends a "single pic back" to the camera + Sends a "singse pic forward" to the camera ++ Sends a "slow motion forward" to the camera

3.3 ARexx Port

The name of the arexx port is ‘LControl_rexx’

So far, the following commands are supported:

‘LC_QUIT’ Quits LControl ‘LC_SHOW’ Opens the LControl window ‘LC_HIDE’ Closese the LControl window

‘LC_ENABLE’

Opens the LControl window and opens the arexx port (if closed)

‘LC_DISABLE’

Closes the LControl window and the arexx port

‘LC_PLAY’
‘LC_STOP’
‘LC_PAUSE’
‘LC_FREW’
‘LC_FFOR’
‘LC_PREW’
‘LC_PFOR’
‘LC_PLUS’
‘LC_MINUS’
‘LC_SLOW’
‘LC_RSLOW’

‘LC_GETCOUNTER’

the current counter will be returned in RESULT

‘LC_RESET’

resets the counter of the camera

‘LC_SWAP’ swaps to the other camera

‘LC_ZOOMIN’

Zooms in one step.

‘LC_ZOOMOUT’

Zooms out one step.

‘LC_RECORD’

Sends Record to the camera.

3.4 Tooltypes

The following tooltypes are recognized:

CX POPUP= "YES" oder "NO"

YES Pop window up on startup

NO Start hidden

CX POPKEY= Hotkey for LControl E.g.: lshift f10 ↦ left shift-key + F10

          Default: lcommand esc.

LEFT=xxx Leftedge of the window

TOP=yyy Topedge of the window PUBSCREEN=name

If specified, LControl tries to open its gui on the pbulicscreen named ‘name’.

Appendix A Support & Bugreports

A.1 Support

Latest software revisions can be found in JESOLO-BBS

JESOLO-BBS, BBS of the "Amiga Freunde Ries" 24H online 2400-19200 Baud, 8N1 Tel.: +49 906 28851

Look in: ‘/PD_POOL/PD-Anwendungen/ANW-Sonstiges/’

You may also send a disk and stamps to one of the following addresses:

Michael Watzl               or      Jürgen Frank
Konradstr. 11                       Wittelsbacherweg 7
86609 Donauwörth                    86609 Donauwörth
Tel.: +49 906 5834                  Tel.: +49 906 1057
Fax : +49 906 28851
Germany                             Germany

A.2 Bugreports

Bugreports or suggestions are welcome and should be sendt to one of the addresses above mentioned.

Please include a brief description of your computer and its configuration — this will help us to reproduce the bugs.

Appendix B Registering

REGISTRATION FORM
FIRM:       ___________________________________________
NAME:       ___________________________________________
STREET:     ___________________________________________
ZIP-CODE:   ______
CITY:       ___________________________________________
COUNTRY:    ___________________________________________
TEL:        _____________
I want to become registered user of LControl, therefore I
o enclose a cheque
  (europe: euro cheque
   others: drawn on a german bank)
o transfer payment to
    Raiffeisen-Volksbank Donauwoerth eG
    86609 Donauwoerth
    Acc:: 4185455   Bank code: 722 901 00
   Comment"LCONTROL REGISTRIERUNG"
Please send me:
o Latest version + scematics for the hardware
  für $12 / DM 20
o Latest version + contruction kit
  für $18 / DM 30
o Latest version + complete Hardware
  für $25 / DM 40
Optional: [for the construction kit & complete hardware]
o Two camera set [allows to connect two cameras and
  switch between them]
  plus $3 / DM 5

Appendix B: Registering 8

         _______________________________________________
            locality, date                            sign

Appendix C IR-Master-Support

Taken from "IR-Master.guide":

"The IR-Master project consists of two software components and a control hardware, which is connected to the joystick port of your Amiga.

At first you start editing your virtual remote control with the IR-Editor… This one is saved to disk then. Afterwards you activate the IR-Runner…

Your edited remote control opens as a window on the workbench. With this window you can interactively control your device[s], you can use the IR-Runner’s ARexx-port for programming your devices… Finally you can send infra red codes from within other applications or edit scripts which will automatically control e.g. your video recorder.

Since version 2.5 of the IR-Master two new features were introduced: You can directly send commands to LControl and Rapid. This allows you to control video cameras supported with an LANC/Control-Linterface and video recorders supported with an

GSE-Rapid-interface. . ."

Appendix D History

V1.00       first Release
V1.01       - arexx-command: GETSTATUS added
            - fixed some minor bugs
V1.02       - Support for 2 cameras added [Switch]
            - minor bugfixes
V1.05       - Zoom in/Zoom out added
            - Record added
To Do:
------
    - localisation

Index 11

Index

-

-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

– . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

+

+ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ++ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

A

About... . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 ARexx Port. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Authors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

B

Bugreports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

C

Control-L . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 CX POPKEY= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 CX POPUP= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

F

FFOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Fix position. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 FREW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

H

Hardware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Hide . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 How to use . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

I

Infra Red . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 IR-Master-Support. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

K

Kickstart . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

L

LANC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 LC DISABLE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC ENABLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC FFOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC FREW. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC GETCOUNTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC HIDE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC MINUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC PAUSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC PFOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC PLAY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC PLUS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC PREW. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC QUIT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC RECORD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC RESET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC RSLOW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC SHOW. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC SLOW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC STOP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC SWAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC ZOOMIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC ZOOMOUT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LEFT=. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

M

Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

P

Parallel port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Pause. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 PFOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Play . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 PREW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Prices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Project . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 PUBSCREEN=. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

Index 12

Q

Quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

R

Record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Registering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Reset counter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

S

Sony. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Special . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Stop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Swap cameras. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

T

Tooltypes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 TOP= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

U

Update . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

V

Videocamera. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

W

What is LControl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Z

Zoom in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Zoom out . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

Table of Contents

1 What is LControl. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 How To Use. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3.1 Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.1.1 Project . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.1.2 Special. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3.2 Gadgets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.3 ARexx Port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.4 Tooltypes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

Appendix A Support & Bugreports. . . . . . . . . . . . . . . . . . . 6

A.1 Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 A.2 Bugreports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

Appendix B Registering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Appendix C IR-Master-Support . . . . . . . . . . . . . . . . . . . . . . 9 Appendix D History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11